A total of $3900 was invested, part of it at 8% interest and the remainder at 13%. If the total yearly interest amounted to $395, how much was invested at each rate?
Answer by mananth(16949)

Investment I 8.00% per annum ---x
Investment II 13.00% per annum ---y

x + y= 3900 ------------------------1
8.00% x + 13.00% y = $395.00
Multiply by 100
8 x + 13 y = $39,500.00

Multiply (1) by -8
we get

-8 x -8 y = -31200.00

Add this to (2)

0 x +5 y = $8,300.00

divide by 5

y = $1,660.00
Balance $2,240.00
CHECK
$2,240.00 --------- 8.00% ------- $179.20
$1,660.00 ------- 13.00% ------- $215.80
Total ------- $395.00